OpenStudy (anonymous):

The cost of renting a van for one day includes a flat rental fee plus a charge for each mile the van is driven while it is rented. A van that is driven 107 miles costs $97.15. A van that is driven 127 miles costs $106.15. What is the flat rental fee? help please! PLEASEEE! :p

OpenStudy (radar):

The flat rate rental fee is $49.00 Do you want to see the steps taken to obtain this solution?

OpenStudy (anonymous):

yes please (: and thank you sooo much for helping me!! i really thankn you!

OpenStudy (radar):

Let F be the flat rate. Let x be the mileage rate. The two equations for the given information would be: F+107x=97.15 and F+127x=106.15 Do you follow this line of reasoning?

OpenStudy (anonymous):

YES! (: thank YOUUUUUUUUUUUUU so much!

OpenStudy (radar):

Can you take it from there, or do you want me to continue?

OpenStudy (anonymous):

if you wannna continue you can, but you dont have to (:

OpenStudy (radar):

I will continue, you may do it different, but I did the following. Take each equation and isolate F. F=97.15-107x F=106.15-127x we know that F=F so we can write: 97.15-107x=106.15-127x Consolidating terms using algebra -107x+127x=106.15-97.15 20x=9.00 x=.45 or 45 cents per mile F=97.15-(107 * .45) F=97.15-48.15 F=$49.00 Did you get the same?

OpenStudy (anonymous):

yes! thank you SO much (:

OpenStudy (radar):

No problem, good day.